OpenStudy (anonymous):

evaluate x-4/2 when x=-1.6

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Is this it? \(\dfrac{x-4}{2}\) Or is it this? \(x - \dfrac{4}{2}\)

OpenStudy (anonymous):

top

OpenStudy (anonymous):

Okay, so let's plug in -1.6 \(\dfrac{x - 4}{2}\) \(\dfrac{-1.6 - 4}{2}\) Subtract -1.6 - 4, what do you get?

OpenStudy (legends):



OpenStudy (anonymous):

-5.6

OpenStudy (legends):

@iGreen. you can do it like this right?

OpenStudy (anonymous):

not hure

OpenStudy (anonymous):

Yes, now divide -5.6 / 2

OpenStudy (anonymous):

No, you can't..

OpenStudy (anonymous):

You have to do the numerator first..then divide by 2.

OpenStudy (anonymous):

-2.8

OpenStudy (anonymous):

igreen i got 2.8